OpenStudy (anonymous):

find the polynomial function that has the given zeros: 3, 2- the square root of 3, 2+ the square root of 3

OpenStudy (campbell_st):

well it will be \[P(x) = (x -3)(x - (2 - \sqrt{3})(x -(2 + \sqrt{3})\]

OpenStudy (anonymous):

what do I do after is what I don't get

OpenStudy (campbell_st):

well distribute would be my thought

OpenStudy (campbell_st):

looking at the last 2 factors you will get \[x^2 - x(2 - \sqrt{3} )-x(2 + \sqrt{3}) + (2 - \sqrt{3}) \times (2 + \sqrt{3})) \] when you get this quadratic multiply it by (x - 3)